### v1

@@ 3.3 The Sequenced aTAM has No Universal Assembly Kit
Some additional definitions will be useful at this point as we leverage Lemma 3.9 to show that deterministic growth of certain shapes with $M$ protrusions from an initial subshape requires at least $\frac{M}{4}$ unique tiles. Let an open face be a face of some non-empty tile in a configuration that neighbours an empty face, and glue types belonging to some open face are similarly known as open glue types. Let $\mathbf{E(A)}$ return the oriented shape of a configuration.
Theorem 3.11.
@@
    For any $i,j\in{1,..,M}$, $S_{i}\not\simeq S_{j}$ for any $i\neq j$. That is, each $S_{i}$ is distinct up to rotations and translations.
@@
    For any $i\in{1,..,M}$, $S_{i}$ has exactly one coordinate that neighbours a coordinate in $S_{0}$, and $S_{0}$ has exactly one coordinate that neighbours this coordinate. We denote by $\dot{z}_{i}$ the coordinate in $S_{i}$ and by $\ddot{z}_{i}$ the neighbouring coordinate in $S_{0}$ (neighbouring faces are similarly labelled ($\dot{z}_{i}$, $\dot{k}_{i}$) and ($\ddot{z}_{i}$, $\ddot{k}_{i}$)).
@@
Consider a sequenced aTAM instance $\mathcal{P}=(A_{0},Q,g)$ such that $E(A_{0})=S_{0}$ and the range of $\mathbf{g}$ is restricted to $\mathbf{g\geq 0}$. If assembly is deterministic and the terminal configuration $A_{\omega}$ is such that $E(A_{\omega})=S$ as defined above, then the number of unique tile types required in $A_{0}$ grows with at least $\frac{M}{4}$.
Proof 3.12.
Condition 4 and determinism imply that $\mathcal{P}$ is oriented-deterministic. Assume $A_{0}$ has fewer than M distinct open glue types, so some $S_{i}$and $S_{j}$ must be anchored at glues of the same type. Then, Theorem 3.9 implies one of the following must be true:
@@
    An overlap of configurations occurs when transforming the configurations neighbouring the two faces $(\ddot{z}_{i},\ddot{k}_{i})$ and $(\ddot{z}_{j},\ddot{k}_{j})$ (in the language of Theorem 3.9, $A_{\omega}^{\prime}=\infty$ and is thus not $g$-valid).
Each of these possibilities leads to a contradiction. The first possibility directly contradicts condition 1. The second possibility results in some point of $S_{i}$ (or $S_{j}$) neighbouring some additional point in $S$ outside of $S_{i}$ (or $S_{j}$), contradicting condition 2 or 3. Hence, by contradiction, $A_{0}$ must have $M$ or more open faces. As each tile type has at most $4$ unique faces, the number of unique tile types in $A_{0}$ must grow with at least $\frac{M}{4}$.
The bolded restriction on $g$ in Theorem 3.11 excludes the possibility of repulsive interactions. Thus we do not have to consider situations where a tile is blocked by some repulsive interaction, rather than overlapping. We have chosen to include this simplified form of the theorem in this text as it provides a better intuition for our proof. However, with a few additional assumptions, we show how this restriction can be lifted in Theorem A.28, included in Appendix A, and the remainder of our results are consistent with this more general form.
We now construct a class of target shapes such that shapes obeying the assumptions in Theorem 3.11 (and more strongly, those in Theorem A.28) cannot be avoided in the assembly of these targets. To do so, we define a branching point as any coordinate of some oriented shape $S$ with three or more neighbours in $S$, while a corner is a coordinate in $S$ with two neighbours in $S$, such that the corner and its two neighbours do not form a straight line. Then, let a straight line segment be a line starting at some branching point or corner, and ending at the next branching point/corner. The distance between two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$ is taken to be $|x_{2}-x_{1}|+|y_{2}-y_{1}|$. We can now begin describing the fundamental components of our constructed target shape.
@@
Note that features 3 and 4 are not necessary for Theorem 3.11, but are necessary for the target shape we are constructing to obey the assumptions of Theorem A.28. We use these oriented shapes to build larger oriented shapes with the aim of creating a class of shapes that fulfill the assumptions laid out in Theorem 3.11 (as well as Theorem A.28). We call these larger oriented shapes treeangles, and we define them as follows.
@@
Proof 3.16.
For simplicity, consider a specific oriented shape $S\in\overline{S}$ and consider the graph $G_{S}(S)$. Let $R_{V}$, $I$ and $\tilde{R}_{V}$ be defined as in Definition 3.14. Then, the subshape defined by any path $P$ from $R_{V}$ or $I$ to a terminal branching point of $\tilde{R}_{V}$, excluding the terminal branching point, must have at least $V$ external neighbours. Consider then a shape $S^{\prime}$ with $P$ as a subshape, such that $S^{\prime}$ does not contain any terminal branching points of $R_{V}$. Without loss of generality, starting assembly at any coordinate of $I$ or $R_{V}^{\prime}$, any complete trajectory $\Psi$ of $\mathcal{P}$ that assembles $S$ must admit an element $A^{\prime}$ such that $E(A^{\prime})=S^{\prime}$ for some $S^{\prime}$ (an analogous argument applies for trajectories starting at a coordinate of $R_{V}$), since there must exist some time point where the first terminal branching point is incorporated onto $S$. Furthermore, $S^{\prime}$ can have no fewer than $V$ external neighbours, because the inclusion of further branching points of $R_{V}$ can only ever increase the number of external neighbours of $S^{\prime}$, and since $S^{\prime}$ does not contain a terminal branching point of $R_{V}$, then the number of external neighbours cannot be decreased by the addition of more coordinates onto $S^{\prime}$. By the construction rules of the treeangle, a partition of $S$ as $S=\bigcup_{i=0}^{M}S_{i}$ following Theorem 3.11 (as well as Theorem A.28 in the Appendix) is then possible with $S_{0}=S^{\prime}$ and $M\geq V$, and the remaining rules are guaranteed by our treeangle construction. Letting $Q^{\prime}$ be the subsequence of $Q$ after the assembly time in which $A^{\prime}$ appears, the sequenced aTAM instance $(A^{\prime},Q^{\prime},g)$ requires at least $\frac{V}{4}$ unique tile types in $A^{\prime}$ to assemble $\overline{S}$ deterministically. Since all non-empty tile types in $A^{\prime}$ must appear in $Q$, $Q$ requires at least $\frac{V}{4}$ unique tile types, completing the proof.
Theorem 3.17.
@@
We only need to show that given any positive integer $V$, some treeangle $\overline{S}$ of order $V$ exists. The number of walks in any $\mathbb{W}_{L,L,L,L}$ increases monotonically with $L$. Each $\mathbb{W}_{L,L,L,L}$ is non-empty as a simple straight line fulfilling length constraints will be in this set. Increasing $L$ monotonically increases the number of shapes in $\mathbb{W}_{L,L,L,L}$, as we can simply extend the walk in the canvas with another straight segment or some number of curved walks. Hence, there is always some $L$ that will provide sufficient walks to generate a treeangle $\overline{S}$ of arbitrary order $V$. This argument extends to other sets $\mathbb{W}_{d,c,c^{\prime},d^{\prime}}$. The proof then follows from Lemma 3.15.
3.4 Comparing the Backboned and Sequenced aTAM: The Role of Geometric Constraints
The loss of geometric constraint (that additional tiles must neighbour immediately preceding tiles) when moving from the backboned aTAM to the sequenced aTAM has two effects. Firstly, the shape space of the sequenced aTAM is much larger than the backboned aTAM. The backboned aTAM’s shape space is equal to the set of self-avoiding walks, which grows with shape size $N$ as $O(2.638^{N})$ from ([5]). By contrast, the shape space of the sequenced aTAM is equal to the shape space of base temperature-1 aTAM (counting only finite shapes), since a sequence $Q$ can always be constructed by identifying the added tile for any assembly sequence of the temperature-1 base aTAM, assuming the latter is single-tile seeded, allows tile rotations and bans specific mismatches. This space, the set of polyominoes, scales as $O(4.5252^{N})$ ([3]).
We might argue that the loss a universal assembly kit for the sequenced aTAM is simply due to this increase in its shape space. However, even though the shape space of the sequenced aTAM scales faster with $N$ than the shape space of the backboned aTAM, that growth is still exponential. The set of sequences of a given length drawing from a finite $k$-size alphabet similarly scales exponentially as $O(k^{N})$. Hence, combinatorial arguments alone do not preclude the existence of some fixed tile type set with $k\geq 5$ that covers the shape space of the sequenced aTAM.
@@ 3.4 Comparing the Backboned and Sequenced aTAM: The Role of Geometric Constraints
The second factor at play is that the restriction in tile placement conveys some advantages in allowing the backboned aTAM to assemble a complex shape with a small number of tiles. Consider, for example, the configuration shown in Figure 13 (a); the sequence of tiles shown and interaction function in Theorem 3.6 will deterministically produce the given shape with the backboned aTAM, starting from an empty configuration. By contrast, using the same model inputs $(A_{empty},Q,g)$ but employing the placement rules of the sequenced aTAM, assembly is not deterministic. The shape in Figure 13 (b) can form if the first blue tile binds to a face on the far side of the growing assembly.
@@
Can the geometric constraint of the backboned aTAM be generalized such that the ability to assemble all shapes in the shape space of the sequenced aTAM? Klarner showed in ([15]) that any polyomino can be uniquely defined by labelling its squares iteratively, starting from neighbours of the squares with the current lowest label (Figure 14.a). It may thus be possible to define a ‘queue-like’ assembly scheme that operates based on this principle, adding future tiles preferentially to existing tiles that were filled earliest in the assembly process (Figure 14.b), although we have not proved this. If so, by returning a sense of geometrical constraint to the sequenced aTAM, a finite universal assembly kit would be recovered (interestingly, this approach requires endowing the assembly scheme with an unbounded memory of the order of added tiles). From this perspective, the loss of a geometrical constraint translates to a loss of shape space-covering power.
Figure 14: (a) Illustration of Klarner’s rules for enumerating squares in a polyomino. Labelling begins at the lowest square in the leftmost column. The next square is selected by considering neighbours of the lowest labelled square, and going clockwise from the face connecting the existing square to the original polyomino (or from the north facing face for the first square; examples in purple and red provided in the figure). Any polyomino can be uniquely enumerated in this way. (b) A possible labelling scheme under an assembly scheme where tiles are added following a queue, preferentially adding tiles to neighbouring tiles added earlier during assembly. Glue $x$ attracts $\underline{x}$, and blank faces are assigned a neutral 00 glue.
@@
If the geometrical restrictions of the backbone appear to provide an advantage, it is natural to ask whether a finite assembly kit exists that allows the sequenced aTAM to deterministically assemble any shape in the shape space of the backboned aTAM. If we allow for neutral faces, identifying such a kit is trivial, as our construction in Theorem 3.2 would still satisfy the sequenced aTAM. However, as we have argued, assembly without neutral interactions is conceptually more interesting. Clearly, the construction in Theorem 3.6 does not work for the sequenced aTAM. When attempting to produce rectangles of m rows and n columns (as in Section 3.2) with the sequenced aTAM $\mathcal{P}_{m,n}=(A_{empty},Q_{m,n},g)$ using a sequence $Q_{m,n}$ and interaction function $g$ derived from the constructive proof of Theorem 3.6, $\mathcal{P}_{m,n}$ fails to deterministically produce the desired rectangle for any $n>2$ (except when $(m,n)=(2,3)$). However, we have thus far been unable to prove or disprove the existence of a finite tile type set for the sequenced aTAM that can deterministically assemble any Hamiltonian path shape without relying on neutral interactions. We now present a partial result towards such a proof.
@@
We propose a scheme for assembling any enclosed shape in the sequenced aTAM without relying on neutral interactions. We define an enclosed shape as any shape with a well defined Hamiltonian-cycle boundary, that additionally also contains all points within this boundary. Additionally, we also require that the Hamiltonian cycle boundary has no coordinate that neighbours more than two coordinates within the boundary.
Theorem 3.19.
@@
There exists a finite tile type set with associated interaction function $g$ such that, for any enclosed shape, a sequenced aTAM instance $(A_{empty},Q,g)$ with elements Of $Q$ drawn from this tile set that deterministically assembles this shape.
Proof 3.20.
@@
[…]
Then, consider a the subset of the interacting directed tiles from Theorem 3.6 that only contain $3$ as a non-backbone glue type. Any enclosed shape can be constructed by drawing the boundary with interacting directed tiles, with the non-backbone glue 3 facing inwards towards the shape interior. Finally, the interaction between the first and last added boundary tile is encoded by $(5,5)$. Then, the remainder of the sequence $Q$ is filled with the tile type $\{4,4,4,4\}$ to fill in the shape (Figure 15).
Figure 15: Scheme for assigning a sequence $Q$ to assemble an enclosed shape. First, an outer boundary is assembled using interacting directed tiles (blue). Then, the interior of the shape is filled (yellow).
We end by noting that, just as tile sequences that deterministically assemble a shape in the backboned aTAM may fail if used in the corresponding sequenced aTAM instance, sequenced aTAM instances cannot generally be directly converted into a backboned aTAM instance that deterministically assembles a shape. For instance, using the scheme in Theorem 3.19 my fail for a backboned aTAM instance. For a large enough shape, the trajectory may take an arbitrary route inside the interior yellow portion of the shape instead of filling it completely.

### v2

@@ 3.3 The Sequenced aTAM has No Universal Assembly Kit
Some additional definitions will be useful at this point as we leverage Lemma 2 to show that deterministic growth of certain shapes with $M$ protrusions from an initial subshape requires at least $\frac{M}{4}$ unique tiles. Let an open face be a face of some non-empty tile in a configuration that neighbours an empty face, and glue types belonging to some open face are similarly known as open glue types. Let $\mathbf{E(A)}$ return the oriented shape of a configuration.
Theorem 3.
@@
    For any $i,j\in{1,..,M}$, $S_{i}\not\simeq S_{j}$ for any $i\neq j$. That is, each $S_{i}$ is distinct including rotations and translations.
@@
    For any $i\in{1,..,M}$, $S_{i}$ has exactly one coordinate that neighbours a coordinate in $S_{0}$, and $S_{0}$ has exactly one coordinate that neighbours this coordinate. We denote by $\dot{z}_{i}$ the coordinate in $S_{i}$ and by $\ddot{z}_{i}$ the neighbouring coordinate in $S_{0}$ (neighbouring faces are similarly labeled ($\dot{z}_{i}$, $\dot{k}_{i}$) and ($\ddot{z}_{i}$, $\ddot{k}_{i}$)).
@@
Consider a sequenced aTAM instance $\mathcal{P}=(A_{0},Q,g)$ such that $E(A_{0})=S_{0}$ and the range of $\mathbf{g}$ is restricted to $\mathbf{g\geq 0}$. If assembly is deterministic and the terminal configuration $A_{\zeta}$ is such that $E(A_{\zeta})=S$ as defined above, then the number of unique tile types required in $A_{0}$ grows with at least $\frac{M}{4}$.
Proof.
Condition 4 and determinism imply that $\mathcal{P}$ is oriented-deterministic. Assume $A_{0}$ has fewer than M distinct open glue types, so some $S_{i}$ and $S_{j}$ must be anchored at glues of the same type. Then, Theorem 2 implies one of the following must be true:
@@
    An overlap of configurations occurs when transforming the configurations neighbouring the two faces $(\ddot{z}_{i},\ddot{k}_{i})$ and $(\ddot{z}_{j},\ddot{k}_{j})$ (in the language of Theorem 2, $A_{\zeta}^{\prime}=\infty$ and is thus not $g$-valid).
Each of these possibilities leads to a contradiction. The first possibility directly contradicts condition 1. The second possibility results in some point of $S_{i}$ (or $S_{j}$) neighbouring some additional point in $S$ outside of $S_{i}$ (or $S_{j}$), contradicting condition 2 or 3. Hence, by contradiction, $A_{0}$ must have $M$ or more open faces. As each tile type has at most $4$ unique faces, the number of unique tile types in $A_{0}$ must grow with at least $\frac{M}{4}$. ∎
The bolded restriction on $g$ in Theorem 3 excludes the possibility of repulsive interactions. Thus we do not have to consider situations where a tile is blocked by some repulsive interaction, rather than overlapping. We have chosen to include this simplified form of the theorem in this text as it provides a better intuition for our proof. However, with a few additional assumptions, we show how this restriction can be lifted in Theorem 6, included in Appendix A, and the remainder of our results are consistent with this more general form.
We now construct a class of target shapes such that shapes obeying the assumptions in Theorem 3 (and more strongly, those in Theorem 6) cannot be avoided in the assembly of these targets. To do so, we define a branching point as any coordinate of some oriented shape $S$ with three or more neighbours in $S$, while a corner is a coordinate in $S$ with two neighbours in $S$, such that the corner and its two neighbours do not form a straight line. Then, let a straight line segment be a line starting at some branching point or corner, and ending at the next branching point/corner. The distance between two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$ is taken to be $|x_{2}-x_{1}|+|y_{2}-y_{1}|$. We can now begin describing the fundamental components of our constructed target shape.
@@
Note that features 3 and 4 are not necessary for Theorem 3, but are necessary for the target shape we are constructing to obey the assumptions of Theorem 6. We use these oriented shapes to build larger oriented shapes with the aim of creating a class of shapes that fulfill the assumptions laid out in Theorem 3 (as well as Theorem 6). We call these larger oriented shapes treeangles, and we define them as follows.
@@
Proof.
For simplicity, consider a specific oriented shape $S\in\overline{S}$ and consider the graph $G_{S}(S)$. Let $R_{V}$, $I$ and $\tilde{R}_{V}$ be defined as in Definition 6. Then, the subshape defined by any path $P$ from $R_{V}$ or $I$ to a terminal branching point of $\tilde{R}_{V}$, excluding the terminal branching point, must have at least $V$ external neighbours. Consider then a shape $S^{\prime}$ with $P$ as a subshape, such that $S^{\prime}$ does not contain any terminal branching points of $R_{V}$. Without loss of generality, starting assembly at any coordinate of $I$ or $R_{V}^{\prime}$, any complete trajectory $\Psi$ of $\mathcal{P}$ that assembles $S$ must admit an element $A^{\prime}$ such that $E(A^{\prime})=S^{\prime}$ for some $S^{\prime}$ (an analogous argument applies for trajectories starting at a coordinate of $R_{V}$), since there must exist some time point where the first terminal branching point is incorporated onto $S$. Furthermore, $S^{\prime}$ can have no fewer than $V$ external neighbours, because the inclusion of further branching points of $R_{V}$ can only ever increase the number of external neighbours of $S^{\prime}$, and since $S^{\prime}$ does not contain a terminal branching point of $R_{V}$, then the number of external neighbours cannot be decreased by the addition of more coordinates onto $S^{\prime}$. By the construction rules of the treeangle, a partition of $S$ as $S=\bigcup_{i=0}^{M}S_{i}$ following Theorem 3 (as well as Theorem 6 in the Appendix) is then possible with $S_{0}=S^{\prime}$ and $M\geq V$, and the remaining rules are guaranteed by our treeangle construction. Letting $Q^{\prime}$ be the subsequence of $Q$ after the assembly time in which $A^{\prime}$ appears, the sequenced aTAM instance $(A^{\prime},Q^{\prime},g)$ requires at least $\frac{V}{4}$ unique tile types in $A^{\prime}$ to assemble $\overline{S}$ deterministically. Since all non-empty tile types in $A^{\prime}$ must appear in $Q$, $Q$ requires at least $\frac{V}{4}$ unique tile types, completing the proof. ∎
Theorem 4.
@@
We only need to show that given any positive integer $V$, some treeangle $\overline{S}$ of order $V$ exists. The number of walks in any $\mathbb{W}_{L,L,L,L}$ increases monotonically with $L$. Each $\mathbb{W}_{L,L,L,L}$ is non-empty as a simple straight line fulfilling length constraints will be in this set. Increasing $L$ monotonically increases the number of shapes in $\mathbb{W}_{L,L,L,L}$, as we can simply extend the walk in the canvas with another straight segment or some number of curved walks. Hence, there is always some $L$ that will provide sufficient walks to generate a treeangle $\overline{S}$ of arbitrary order $V$. This argument extends to other sets $\mathbb{W}_{d,c,c^{\prime},d^{\prime}}$. The proof then follows from Lemma 3. ∎
3.4 Geometric Constraints, Shape Space Size and Kolmogorov Complexity
We now frame our results in the context of the shape space and worst-case Kolmogorov complexity scaling of our considered assembly maps. Sequences of length $N$, drawing from an alphabet of size $M$, have $M^{N}$ unique variants, and hence can in principle deterministically encode up to $M^{N}$ shapes. Similarly, an arbitrary sequence of length $N$ requires $N\log{M}$ bits, and hence the “program” defined by a backboned or sequenced aTAM exceeds $N\log{M}$ in length, scaling as $O(N)$ if $M$ is fixed.
The backboned aTAM’s shape space is equal to the set of self-avoiding walks, which grows with shape size $N$ as $O(2.638^{N})$ from (?). Further, as self-avoiding walks can be encoded by left-right-forward moves, the scaling in worst-case Kolmogorov complexity is upper bounded by $O(N)$. Consistent with Theorems 1 and 2, these scaling results do not preclude a universal assembly kit of size $M\geq 3$. Indeed, the backboned aTAM without restrictions gets very close to a universal assembly kit with $M=4$, perhaps unsurprisingly as it simply routes the underlying hamiltonian path. Constraining the assembly mechanism so that all interactions must be attractive does not prevent the existence of a universal assembly kit, but our upper bound on the kit size – over $200$ tiles – suggests that the information in the sequence is used much less efficiently. Although the true minimal universal assembly kit size may be smaller than 200 tiles, it will likely share the same two sources of inefficiency: firstly, many sequences will form equivalent structures; and secondly, others will not deterministically form a well-defined structure.
It is interesting to note the interplay between these two sources of inefficiency. Imposing more demanding rules on the assembly mechanism forces more tiles to be used to specify the same set of backbone routings, in order to avoid non-deterministic or incomplete assemblies. In turn, this larger set of tiles leads to a redundancy, whereby many sequences deterministically assemble the same structure.
The shape space of the sequenced aTAM is equal to the set of free polyominoes (polyominoes equivalent under rotation), with a scaling upper bound given by $O(4.5252^{N})$ (?) (this upper bound is presented for fixed polyominoes, but note the numbers of fixed and free polyominoes are related by a constant factor in the large $N$ limit (?)). We now also argue that the scaling of the worst case Kolmogorov complexity of shapes in the shape space of sequenced aTAM is upper bounded by $O(N)$.
While to our knowledge this argument for the $O(N)$ scaling has not been explicitly made in literature, it follows trivially from Klarner’s algorithm for polyomino enumeration (?). A brief review of Klarner’s algorithm is provided as Algorithm 1. Starting at the lowest tile of the leftmost column of a fixed polyomino $S$, Klarner’s algorithm iteratively connects new tiles to neighbouring “parent” tiles, creating a queue of tiles as it goes. The tiles in the queue are successively tested for new neighbours. Moreover, each tile is labelled by going clockwise around the tile, starting from (but excluding) the given tile’s parent. The label is a 3-bit description, storing a value ‘$1$’ if a tile’s neighbor is present in $S$, and ‘00’ otherwise, for each neighbour (Figure 14).
1
2Function getNeighbors($S,\mathrm{current},\mathrm{parent},\&\mathrm{tileQueue},\&\mathrm{parentMap}$):
3       triple $\leftarrow(0,0,0)$
4       for $i\leftarrow 1$ to $3$ do
5             neighbor $\leftarrow$ neighbor $i$ of current, counting clockwise with parent at $i=0$
6             if $\mathrm{neighbor}\in S\mathrm{\;and\;}\mathrm{parentMap}(\mathrm{neighbor})=\mathrm{empty}$ then
7                   triple$[i]\leftarrow 1$
8                   parentMap((((neighbor)$\leftarrow$) ←current
9                   enqueue((((tileQueue,neighbor))))
10
11      return $\mathrm{triple}$
12
13
Input: Fixed polyomino $S$
14 $N\leftarrow$ size$(S)$
15 current $\leftarrow$ lowest tile position of the leftmost column of $S$
16 parent $\leftarrow$ position below current
[…]
18 parentMap $\leftarrow\mathrm{Map}\{\mathrm{current^{\prime}}\to\mathrm{parent^{\prime}}\}$
19 result $\leftarrow$ array of $N$ triples $(0,0,0)$
20
21for $i\leftarrow 1$ to $N$ do
22       result$[i]\leftarrow$ getNeighbors($S,\mathrm{current},\mathrm{parent},\mathrm{tileQueue},\mathrm{parentMap}$)
23       current $\leftarrow$ dequeue((((tileQueue))))
24       parent $\leftarrow$ parent_map((((current))))
25return $\mathrm{result}$
Algorithm 1 Klarner’s algorithm for uniquely labelling polyominoes. A $3N$-bit description of a given input polyomino of size $N$ is produced as output.
Figure 14: Illustration of Klarner’s rules for enumerating squares in a polyomino. Labeling begins at the lowest square in the leftmost column. The next square is selected by considering neighbors of the lowest labeled square, and going clockwise from the face connecting the existing square to the original polyomino (or from the north facing face for the first square; examples in purple and red provided in the figure). A bit description for each tile is then written down, recording tiles which ‘grow’ from that particular tile. For example, tile $2$ has the description $(0,0,1)$, while tile $4$ has the description $(0,1,0)$.
The output of Klarner’s algorithm is therefore a series of bits of length $3N$ (note the first bit is omitted in Klarner’s original argument, leading to $13N-$ 1 bits. For convenience, this bit exists in our implementation but is always ‘0’). Importantly, every polyomino is assigned a unique bit description, and there exists an inverse algorithm to recover a given fixed polyomino from Klarner’s bit description, and the length of this program does not scale with $N$ (Algorithm 2). Hence, the scaling of the worst case Kolmogorov Complexity of a an algorithm for consturcting any given shape of size $N$ is upper bounded by $O(N)$ (note that adding equivalence between rotations at worst adds a constant to the program length).
Once again, these bounds do not preclude a universal assembly kit, this time with $M\geq 5$. For some finite $M\geq 5$ there are more sequences of length $N$ than shapes of size $N$ in the shape space; for some finite $M$ there are more bits stored in a sequence of length $N$ than bits required to write a program that constructs any shape of size $N$ in the shape space. Despite these facts, however, as Theorem 4 shows, no universal assembly kit exists for the sequenced aTAM. The inefficiency of the assembly rules are so strong that, not only is the number of tiles far larger than $M=5$, no finite $M$ is sufficient at all.
Klarner’s algorithm itself gives us some insight into the cause of this inefficiency. If one were able to use a queue, as in Klarner’s algorithm, so that new tiles could only be added adjacent to the tile in a configuration that was at the top of a queue, it would be straightforward to identify a universal assembly kit. This queue-based geometric constraint would effectively be the generalization of the backbone to a process that allows for branching points, rather than requiring a single self-avoiding walk. However, implementing such a queuing system in a molecular asssembly process rather than in silico would be highly non-trivial.
Input: array of bit triples tripleList
1 $N\leftarrow$ size((((tripleList))))
2 $S\leftarrow$ array($(\{(0,0)\})$
3 parentMap $\leftarrow\mathrm{Map}\{\mathrm{current^{\prime}}\to\mathrm{parent^{\prime}}\}$
4 parentMap$((0,0))\leftarrow(0,-1)$
5 for $i\leftarrow 1$ to $N$ do
6       current $\leftarrow S[i]$
7       parent $\leftarrow$ parentMap((((current))))
8       for $j\leftarrow 1$ to $3$ do
9             if $\mathrm{tripleList}[i][j]$ then
10                   neighbor $\leftarrow$ neighbor $j$ of current, clockwise with parent at $j=0$
11                   parentMap(neighbor) $\leftarrow$ current
12                   $S$.pushBack(neighbor)
13
14
15return $S$
Algorithm 2 Klarner’s inverse algorithm for obtaining the coordinates of a given polyomino from its bit description.
3.5 Applying the sequenced aTAM to the shape space of the backboned aTAM
@@
Geometric restrictions in tile placement therefore appear to convey some advantages in allowing the assembly of a complex shape with a small number of tiles. For a concrete example, in addition to the discussion in Section 3.4, consider the configuration shown in Figure 15 (a); the sequence of tiles shown and interaction function in Theorem 2 will deterministically produce the given shape with the backboned aTAM, starting from an empty configuration. By contrast, using the same model inputs $(A_{empty},Q,g)$ but employing the placement rules of the sequenced aTAM, assembly is not deterministic. The shape in Figure 15 (b) can form if the first blue tile binds to a face on the far side of the growing assembly.
@@
If the geometrical restrictions of the backbone appear to provide an advantage, it is natural to ask whether a finite assembly kit exists that allows the sequenced aTAM to deterministically assemble any shape in the shape space of the backboned aTAM. If we allow for non-attractive inter-tile interactions, identifying such a kit is trivial, as our construction in Theorem 1 would still satisfy the sequenced aTAM. However, as we have argued, assembly with only attractive inter-tile interactions is conceptually more interesting. Clearly, the construction in Theorem 2 does not work for the sequenced aTAM. When attempting to produce rectangles of $m$ rows and $n$ columns (as in Section 3.2) with the sequenced aTAM $\mathcal{P}_{m,n}=(A_{empty},Q_{m,n},g)$ using a sequence $Q_{m,n}$ and interaction function $g$ derived from the constructive proof of Theorem 2, $\mathcal{P}_{m,n}$ fails to deterministically produce the desired rectangle for any $n>2$ (except when $(m,n)=(2,3)$). However, we have thus far been unable to prove or disprove the existence of a finite tile type set for the sequenced aTAM that can deterministically assemble any Hamiltonian path shape with only attractive inter-tile interactions. We now present a partial result towards such a proof.
@@
We propose a scheme for assembling any enclosed shape in the sequenced aTAM without relying on non-attractive inter-tile interactions. We define an enclosed shape as any shape with a well defined Hamiltonian-cycle boundary, that additionally also contains all points within this boundary. Additionally, we also require that the Hamiltonian cycle boundary has no coordinate that neighbours more than two coordinates within the boundary.
Figure 16: Scheme for assigning a sequence of tile types to assemble (a) rectangles and (b) bulged rectangles, in the sequenced aTAM with only attractive inter-tile interactions. First, an outer boundary is assembled using interacting directed tiles (blue). Then, the interior of the shape is filled (yellow). Note the bulged rectangle itself is not an enclosed shape, but the same shape filling approach can be applied after adjusting for faces in the bulge.
Theorem 5.
@@
There exists a finite tile type set with associated interaction function $g$ such that, for any enclosed shape, a sequenced aTAM instance $(A_{empty},Q,g)$ with elements of $Q$ drawn from this tile set that deterministically assembles this shape with only attractive inter-tile interactions.
Proof.
@@
$g(\sigma,\sigma^{\prime})=\left\{\begin{array}[]{ll}1,&\mbox{if}\ (\sigma,\sigma^{\prime})=(1,2)\mbox{, }(\sigma,\sigma^{\prime})=(3,4)\mbox{, }(\sigma,\sigma^{\prime})=(4,4)\mbox{ or }(\sigma,\sigma^{\prime})=(5,5)\\ -3,&\mbox{otherwise}.\end{array}\right.$ (3)
Then, consider a the subset of the interacting directed tiles from Theorem 2 that only contain $3$ as a non-backbone glue type. Any enclosed shape can be constructed by drawing the boundary with interacting directed tiles, with the non-backbone glue 3 facing inwards towards the shape interior. Finally, the interaction between the first and last added boundary tile is encoded by $(5,5)$. Then, the remainder of the sequence $Q$ is filled with the tile type $\{4,4,4,4\}$ to fill in the shape (For an example, refer to Figure 16(a)).∎
Figure 17: Upper bounds on the tile complexity, or the number of tiles needed to construct (a) rectangles and (b) bulged rectangles of various row and column numbers using sequenced aTAM, with and without interaction restrictions. There is an increase in the number of tiles needed to assemble both shapes when inter-tile interactions are forced to be attractive, but in general this increase is small.
We now consider upper bounds on the number of tiles required to produce the rectangle and bulged rectangle shapes in Figures 17(a) and 17(b) with unconstrained Sequenced aTAM and Sequenced aTAM with only attractive inter-tile interactions. With the former, it is possible to assemble these shapes in the same way as backboned aTAM, using left-right-forward tiles to follow a hamiltonian path. With the latter, we apply the shape filling approach in Theorem 5, making slight adjustments for the bulge in the bulged rectangle (note the bulged rectangle itself is not an enclosed shape, but the shape filling approach can be applied after this adjustment). Results are provided in Figure 17. For shapes where shape filling is possible, it is quite efficient, only requiring a few more tiles than unconstrained sequenced aTAM.
Moreover, we end by noting that, just as tile sequences that deterministically assemble a shape in the backboned aTAM may fail if used in the corresponding sequenced aTAM instance, sequenced aTAM instances cannot generally be directly converted into a backboned aTAM instance that deterministically assembles a shape. For instance, using the scheme in Theorem 5 may fail for a backboned aTAM instance. For a large enough shape, the trajectory may take an arbitrary route inside the interior yellow portion of the shape instead of filling it completely.